OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 6) and (2, 12)? f(x) = 3(2)x f(x) = 2(3)x f(x) = 3(2)−x f(x) = 2(3)−x

OpenStudy (campbell_st):

just substitute the x values to see if you get the y values 1st point x = 1 and f(1) = 6 so does \[6 = 3 \times 2^1\] is it's true, then use the 2nd point x = 2 and f(x) = 12 and see if it's true again... two trues will give the correct answer

OpenStudy (anonymous):

Is it b? @campbell_st

OpenStudy (campbell_st):

well looking at A 6 = 3 x 2 which is true in B you have \[6 = 2 \times 3^1\] which is also true... so take the 2nd point x = 2 and f(2) = 12 and see which of A or B is true...

OpenStudy (anonymous):

Oh okay is it a? @campbell_st

OpenStudy (campbell_st):

is that a guess or because you calculated an answer

OpenStudy (anonymous):

I calculated @campbell_st

OpenStudy (anonymous):

or at least i tried

OpenStudy (campbell_st):

so you didn't need to ask... you have the answer A. \[12 = 3 \times 2^2\] B. \[12 = 2 \times 3^2\]

OpenStudy (anonymous):

oh ok thank you!